# Education-Oriented Graph Retrieval-Augmented Generation for Learning Path Recommendation

### v1

@@ Abstract.
Learning path recommendation seeks to provide learners with a structured sequence of learning items (e.g., knowledge concepts or exercises) to optimize their learning efficiency. Despite significant efforts in this area, most existing methods primarily rely on prerequisite relationships, which present two major limitations: 1) Many educational datasets do not explicitly provide prerequisite relationships between knowledge concepts, hindering the application of current learning path recommendation methods. 2) Relying solely on prerequisite relationships as the sole knowledge structure can impede learning progress and negatively impact student outcomes. To address these challenges, we propose a novel approach, Discrimination Learning Enhances Learning Path Recommendation (DLELP), which enhances learning path recommendations by incorporating both prerequisite and similarity relationships between knowledge concepts. Specifically, we introduce a knowledge concept structure graph generation module that adaptively constructs knowledge concept structure graphs for different educational datasets, significantly improving the generalizability of learning path recommendation methods. We then propose a Discrimination Learning-driven Reinforcement Learning (DLRL) framework, which mitigates the issue of blocked learning paths, further enhancing the efficacy of learning path recommendations. Finally, we conduct extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides interpretable reasoning for the recommended learning paths.
@@
copyright: noneconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: ccs: Applied computing E-learningccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. The comparison of traditional methods (top) and our proposed DLELP (bottom) is depicted. The top denotes traditional KC prerequisite relationship-based methods where confusion between similar KCs such as $\text{KC}_{3}$, $\text{KC}_{6}$, and $\text{KC}_{7}$ can hinder students’ mastery of $\text{KC}_{3}$ and impede their understanding of subsequent KCs. The bottom illustrates how DLELP enables students to more effectively differentiate between these similar KCs, thereby improving learning outcomes.
With the rapid advancement of online education platforms, the growing demand for personalized learning experiences has led to the increasing significance of learning path recommendation (LPR) (Wang et al., 2023; Zhang et al., 2021; Chen, 2009; Shi et al., 2020). It can adaptively recommend personalized learning paths based on individual learning tempo, enabling them to acquire knowledge and skills systematically and orderly, unlike the ”general education” approach in traditional education (Nabizadeh et al., 2017; Guan et al., 2025; Cheng et al., 2025).
Research on education has shown that the structure of the knowledge concept (KC) significantly influences the LPR (Liu et al., 2019). Existing solutions based on KC structure for LPR can be mainly divided into two parts: KC correlation-based, where personalized learning paths are provided to learners by leveraging the implicit correlations among knowledge concepts, and KC prerequisite relation-based, where personalized learning paths are offered to learners by leveraging the knowledge concepts prerequisite relationships. Among them, KC correlation-based methods, without considering the unambiguous KC structure, fail to precisely determine the customized learning tempo (Chen et al., 2023). In contrast, KC prerequisite relationship-based methods integrate learners’ knowledge mastery levels with the prerequisite relationships of learning materials to provide smooth and efficient learning paths (Liu et al., 2019; Li et al., 2023; Zhang et al., 2024; Li et al., 2024). Despite the significant success of these methods, they often encounter the following two challenges.
Most educational datasets do not explicitly provide prerequisite relationships between knowledge concepts, which hinders the application of existing LPR (Zhang et al., 2024). While few datasets provide corresponding KC prerequisite relationship graphs, these graphs often suffer from sparsity due to the incomplete knowledge concepts or inappropriate relationships in the proposed connections. Additionally, existing educational datasets rely on manually annotated KC prerequisite relationship graphs, which are labor-intensive to construct. Such issues restrict the applicability of existing learning path recommendation methods.
Relying solely on prerequisite relationships between KCs as the single knowledge structure can lead to ”blocked phenomena,” which hinders learning progress (Pan et al., 2024). A major cause is the lack of opportunities for learners to discriminate between similar KCs during the learning process. When confusion arises among these similar KCs, it impedes the mastery of key KC along the learning path, thereby obstructing the acquisition of target KC. According to Gagne’s learning hierarchy (Chinda, 2022), discrimination learning is a critical stage in the learning process, where learners develop the ability to distinguish or discriminate between two KCs (Rohrer, 2012). Figure 1 demonstrates that utilizing similar KCs for discrimination learning can significantly enhance the likelihood of a smoother learning path, improving both learning efficiency and outcomes. Existing traditional learning path recommendation methods lie in the exclusive reliance on sparse KC prerequisite relationships while disregarding KC similarity relationships, which consequently impedes the ability to recommend efficient learning paths that align with learners’ pacing.
To address the aforementioned challenges, this paper proposes a novel method named Discrimination Learning Enhances Learning Path Recommendation (DLELP) to achieve a more satisfactory learning path by considering dual KC structure prerequisite relationships and similarity relationships. Specifically, we propose a KC structure graph generation module to adaptively generate KC structure graphs. This module leverages TextGrad (Yuksekgonul et al., 2024) to mitigate hallucination issues in LLMs, thereby adaptively generating more accurate and densely connected KC structure graphs tailored to different educational datasets. Next, we develop a Discrimination Learning-driven Reinforcement Learning (DLRL) framework based on the generated KC prerequisite relationship graph and similarity relationship graph, which consists of a prerequisite agent and a similarity agent. The prerequisite agent determines whether learning should follow the prerequisite relationships among KCs. The similarity agent identifies easily confused KCs during blocked phenomena and integrates them into the learning path for discrimination learning.
@@
  • We propose a KC structure graph generation module that adaptively generates prerequisite relationship graphs and similarity relationship graphs for KCs based on different educational datasets, significantly enhancing the generalizability of various learning path recommendation methods.
  • We develop a discrimination learning-driven reinforcement learning (DLRL) framework that alleviates the blocked phenomena, further improving the performance of learning path recommendations.
  • We conduct extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides reasons for the recommended learning paths.
@@ 3. Preliminaries
In this section, we discuss the formulation of learning path recommendations and the definition of Graph RAG (Edge et al., 2024).
3.1. Problem Definition
@@
The primary objective of this paper is to solve the knowledge concept structure-based goal-oriented learning path recommendation problem. Suppose we have a set of exercise $E=\{e_{1},e_{2},\cdots,e_{M}\}$ and a set of knowledge concepts $C=\{c_{1},c_{2},\cdots,c_{N}\}$. Normally, for each learner, learning goals are denoted as $G=\{g_{1},g_{2},\cdots\}$, which is a subset of the exercise set $G\subset C$. Taking historical learning records $H=\{h_{0},h_{1},\cdots\,h_{i}\}$ as input and guided by the learning goal, a learning path $P=\{e_{1},e_{2},\cdots,e_{k}\}$ is recommended for learner. Each record $h_{i}=(e,score)$ contains a exercise $e$ and the corresponding performance score.
@@
At the start of each learning session, learner gets an initial score $score_{i}$, where a score of $1$ indicates that the learner has already mastered the learning goal and 00 indicates otherwise. During each learning step, learner tackles an exercise $e_{t}$ and gives a feedback $\text{score}_{t}$. The series of exercises engaged by the learner constitutes the learning path. At the end of the session, a final feedback $\text{score}_{m}$ is provided by learning. To evaluate the effectiveness of the recommended learning path, we calculate the improvement in the learner’s feedback scores between the beginning and the end of the session. The effectiveness of the learning path recommendation, denoted as $E_{p}$ is computed as follows:
(1) $E_{p}=\frac{E_{\text{end}}-E_{\text{start}}}{E_{\text{sup}}-E_{\text{start}}},$
@@
where $E_{\text{sup}}$ represents the maximum achievable score for the learning goals, which corresponds to the total number of learning objectives. Our objective is to maximize the $E_{p}$ by provide an effective learning path.
Figure 2. DLELP Framework Overview. The KC structure graph generation module on the left begins by inputting KC names into TextGrad, which then generates textual descriptions. These descriptions are subsequently used by the EDU-graph RAG to construct KC structure graphs. On the right, the learning path generation module initiates with the P-Agent and S-Agent filtering KCs. Following this, the D-Agent selects exercises of appropriate difficulty. The module then utilizes an environment KT to dynamically update the knowledge state in each session. Finally, the module generates the learning path step by step.
Figure 3. Knowledge Concept Analysis Objective.
3.2. Graph Retrieval-Augmented Generation
Graph RAG (Edge et al., 2024) is an open-source method developed by Microsoft, which is based on global summarization of an LLM-derived knowledge graph. It excels in capturing semantic information, extracting entities and relationships from text, and constructing corresponding knowledge graphs.
The workflow of Graph RAG involves two primary processes: indexing and querying. During the indexing process, the input textual data is divided into chunks. For each text unit, entities, relationships, and claims are identified and extracted. Using this extracted information, a knowledge graph is constructed. The knowledge graph is then hierarchically clustered to generate multiple community-level summaries. In the querying process, the knowledge graph and the generated summaries from the indexing stage are leveraged for inference based on the requirements of specific query tasks. Additionally, Graph RAG’s prompts can be tailored to meet the needs of various tasks across different domains, ensuring its adaptability to a wide range of applications.
To achieve a holistic understanding of the textual descriptions of knowledge concepts and generate accurate KC structure graphs, this paper introduces an Edu-Graph RAG method based on Graph RAG. This approach effectively alleviates the issue of insufficient textual content in knowledge concept descriptions within datasets and mitigates hallucination problems in LLMs, enabling the generation of prerequisite relation and similarity relation graphs for knowledge concept global summarization.
3.3. Framework Overview
Figure 2 presents an overview of DLELP framework, which consists of two primary modules. The left KC structure graph generation module addresses the limitation that most educational datasets do not explicitly provide prerequisite relationships and similarity relationships between KCs. This is achieved by leveraging graph retrieval-augmented generation method (Graph RAG) tailored to the education domain. By analyaing the textual descriptions of KCs, the module identifies prerequisite and similarity relationships to construct KC structure graphs. The right DLRL framework operates interactively during each session. It employs relationship agents that integrate both prerequisite relationships and similarity relationships among KCs to dynamically generate the final learning path.
4. KC Structure Graph Generation
In this section, we provide a concise overview of the KC structure graph generation module. First, we incorporate the TextGrad (Yuksekgonul et al., 2024) technique to analyze KC names, facilitating a more comprehensive acquisition of explanatory texts. Building on this, we design an EDU-Graph RAG approach to automatically generate KC structure graphs. By leveraging the generated textual descriptions, this module effectively captures both prerequisite and similarity relationships, constructing KC structure graphs that serve as a critical foundation for recommending appropriate learning paths.
4.1. KC Textual Analysis and Generation
To effectively capture the rich semantic relationships between KCs, we first parse them to generate detailed explanatory texts as a foundation for further analysis. Given that existing datasets for learning path recommendation lack detailed explanatory texts for KCs, we employ LLMs (Zhao et al., 2023) to generate these explanations, significantly lowering the effort required for manual verification and explanation generation. While LLMs can quickly retrieve and generate explanations via online searches, they are prone to hallucinations, which can lead to inaccurate explanations of the KCs.
To enhance the accuracy of knowledge concept analysis, we utilize the TextGrad (Yuksekgonul et al., 2024) method to evaluate the initial explanations generated by LLMs and iteratively refine them based on the evaluation results. The TextGrad workflow is represented as follows:

### v2

# GraphRAG-Induced Dual Knowledge Structure Graphs for Personalized Learning Path Recommendation

@@ Abstract.
Learning path recommendation seeks to provide learners with a structured sequence of learning items (e.g., knowledge concepts or exercises) to optimize their learning efficiency. Despite significant efforts in this area, most existing methods primarily rely on prerequisite relationships, which present two major limitations: 1) Requiring prerequisite relationships between knowledge concepts, which are difficult to obtain due to the cost of expert annotation, hindering the application of current learning path recommendation methods. 2) Relying on a single, sequentially dependent knowledge structure based on prerequisite relationships implies that difficulties at any stage can cause learning blockages, which in turn disrupt subsequent learning processes. To address these challenges, we propose a novel approach, GraphRAG-Induced Dual Knowledge Structure Graphs for Personalized Learning Path Recommendation (KnowLP), which enhances learning path recommendations by incorporating both prerequisite and similarity relationships between knowledge concepts. Specifically, we introduce a knowledge concept structure graph generation module EDU-GraphRAG that adaptively constructs knowledge concept structure graphs for different educational datasets, significantly improving the generalizability of learning path recommendation methods. We then propose a Discrimination Learning-driven Reinforcement Learning (DLRL) module, which mitigates the issue of blocked learning paths, further enhancing the efficacy of learning path recommendations. Finally, we conduct extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides interpretable reasoning for the recommended learning paths.
@@
copyright: noneconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: $c$s: Applied computing E-learningccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. The comparison of traditional methods (top) and our proposed KnowLP (bottom) is depicted. The top denotes traditional KC prerequisite relationship-based methods where confusion between similar KCs such as $\text{KC}_{3}$, $\text{KC}_{6}$, and $\text{KC}_{7}$ can hinder students’ mastery of $\text{KC}_{3}$ and impede their understanding of subsequent KCs. The bottom illustrates how KnowLP enables students to more effectively differentiate between these similar KCs, thereby improving learning outcomes.
With the rapid advancement of online education platforms, there is a growing demand for personalized learning experiences tailored to individual learners. In this context, Learning Path Recommendation (LPR) has emerged as a critical task, aiming to construct coherent and customized sequences of learning resources that align with specific educational goals (Wang et al., 2023; Zhang et al., 2021; Chen, 2009; Shi et al., 2020). By organizing content in a pedagogically meaningful order, LPR facilitates systematic and progressive knowledge acquisition, thereby enhancing learning efficiency and effectiveness (Nabizadeh et al., 2017; Guan et al., 2025; Cheng et al., 2025).
Prior research in education has demonstrated that the structure of Knowledge Concepts (KCs) plays a pivotal role in shaping effective learning path recommendations (Liu et al., 2019). Existing LPR approaches that incorporate KC structure can be broadly categorized into two paradigms: (i) correlation-based methods (Chen et al., 2023), which infer personalized learning paths by exploiting implicit semantic or statistical associations among KCs; and (ii) prerequisite-based methods (Li et al., 2024; Zhang et al., 2024), which utilize explicit prerequisite relationships to guide the sequential organization of learning content. Among these, prerequisite-based approaches have gained particular attention due to their strong pedagogical grounding and consistency with human cognitive development, offering more interpretable and cognitively aligned learning trajectories.
Prerequisite relationships among KCs serve as a foundational scaffold for learning path recommendation, enabling stepwise progression toward learning objectives. However, the construction of such relationships often relies on costly expert annotations, thereby limiting the applicability of prerequisite-based methods (Zhang et al., 2024). Even when prerequisite graphs are available, they frequently suffer from sparsity and noise due to missing concepts or inaccurate dependencies. Recent advances in Large Language Models (LLMs) offer a compelling alternative by inferring prerequisite structures from natural language. Yet, their effectiveness remains constrained by the limited availability of high-quality textual descriptions in practice and the inherent risk of generating hallucinated or misleading links.
Moreover, prerequisite relationships often enforce a rigid sequential structure, where difficulties in mastering a single concept can cascade into broader disruptions along the learning path (Pan et al., 2024). This structural inflexibility underscores the need for adaptive mechanisms that enable learners to recover when encountering conceptual obstacles. According to Gagné’s learning hierarchy (Chinda, 2022), discrimination learning is essential for distinguishing between closely related knowledge concepts (KCs) (Rohrer, 2012). Integrating similar KCs into the learning process provides a contrastive scaffold, allowing learners to resolve ambiguities and strengthen conceptual differentiation. As shown in Figure 1, this approach improves learning resilience and promotes smoother progression, even in the presence of incomplete or uncertain prerequisite structures.
To address the aforementioned challenges, we propose KnowLP (Knowledge structure graph retrieval-augmented generation for Learning Path recommendation), a novel framework that jointly incorporates prerequisite and similarity relationships between KCs to construct more effective and adaptive learning paths. Specifically, we introduce a KC structure graph generation module EDU-GraphRAG that adaptively constructs task-specific KC graphs. This module leverages TextGrad (Yuksekgonul et al., 2025) to reduce LLM hallucinations, enabling the generation of accurate and densely connected KC structures tailored to diverse educational datasets. Based on the generated graphs, we design a Discrimination Learning-driven Reinforcement Learning (DLRL) module composed of two cooperative agents: a prerequisite agent and a similarity agent. The prerequisite agent decides whether the learning trajectory should adhere to prerequisite dependencies, while the similarity agent detects conceptually confusable KCs in blocked learning scenarios and integrates them into the learning path to promote discrimination learning.
@@
  • We propose KnowLP, a novel framework for personalized learning path recommendation, which adaptively constructs KC relationships from the target data and effectively generates learning paths through the collaboration of prerequisite, similarity, and difficulty agents.
  • We propose an EDU-GraphRAG module that adaptively generates prerequisite and similarity relationship graphs for KCs on different datasets, with the goal of completing missing or incomplete knowledge structure graphs.
  • We develop a DLRL module that introduces similarity relationships as a fallback when prerequisite relationships become ineffective, thereby alleviating the blocked phenomena.
  • We conduct extensive experiments on three educational datasets, demonstrating that our method not only achieves state-of-the-art performance but also generates more effective longer learning paths.
@@ 2.2. Retrieval-Augmented Generation
Figure 2. KnowLP Framework Overview. The KC structure graph generation module on the left begins by inputting KC names into TextGrad, which then generates textual descriptions. These descriptions are subsequently used by the EDU-GraphRAG to construct KC structure graphs. On the right, the learning path generation module initiates with the P-Agent and S-Agent filtering KCs. Following this, the D-Agent selects exercises of appropriate difficulty. The module then utilizes an environment KT to dynamically update the knowledge state in each session. Finally, the module generates the learning path step by step.
@@ 3. Preliminaries
In this section, we discuss the formulation of learning path recommendations and the definition of GraphRAG.
Problem Definition
@@
The primary objective of this paper is to address the knowledge concept (KC) structure-based, goal-oriented learning path recommendation problem. Suppose we have a set of exercises $E=\{e_{1},e_{2},\cdots,e_{M}\}$ and a set of KCs $C=\{c_{1},c_{2},\cdots,c_{N}\}$. For each learner, learning goals are typically defined as $G=\{g_{1},g_{2},\cdots\}$, where $G\subseteq C$. Each historical record $h=(e,\text{score})$ consists of an exercise $e\in E$ and the corresponding performance score. Given a sequence of such records $H=\{h_{1},h_{2},\cdots,h_{i}\}$, where each $h_{j}=(e_{j},\text{score}_{j})$, and guided by the learning goal $g\in G$, a learning path $P=\{\tilde{e}_{1},\tilde{e}_{2},\cdots,\tilde{e}_{k}\}$ is recommended, where $\tilde{e}_{j}\in E$ denotes a recommended exercise.
@@
To evaluate the effectiveness of the recommended learning path, we calculate the improvement in the learner’s feedback scores between the beginning and the end of the session. The effectiveness of the learning path recommendation, denoted as $E_{p}$ is computed as follows:
(1) $E_{p}=\frac{E_{\mathrm{end}}-E_{\mathrm{start}}}{E_{\mathrm{sup}}-E_{\mathrm{start}}},$
@@
where $E_{\mathrm{sup}}$ represents the maximum achievable score for the learning goals, which corresponds to the total number of learning objectives. Our objective is to maximize the $E_{p}$ by provide an effective learning path.
Graph Retrieval-Augmented Generation
Given a natural language query $Q$ and a source document $D$, GraphRAG (Edge et al., 2024) aims to transform unstructured textual content into a structured graph representation $\phi(D)$, which facilitates accurate and explainable answer generation. To achieve this, the extracted semantic elements, such as entities and relations, are passed to a graph constructor $\mathcal{G}(\cdot,\cdot)$, which aggregates and deduplicates them into coherent local subgraphs. At query time, the model generates the final output $A$ by conditioning an LLM on the query $Q$ and the structured graph index $\phi(D)$:
(2) $A=\mathrm{LLM}(Q,\phi(D))=\mathrm{LLM}(Q,\bigcup_{i=1}^{n}\mathcal{G}\left(\mathcal{E},\mathcal{R}\right)),$
where $\mathcal{E}$ denotes the set of extracted entities and $\mathcal{R}$ denotes the set of corresponding relationships among them.
4. Methodology
In this section, we introduce our approach, KnowLP, shown in Figure 2, which consists of two main components: Knowledge Structure Graph Generation and Learning Path Generation. Knowledge Structure Graph Generation is responsible for constructing high-quality knowledge concept (KC) structure graphs by adaptively discovering both prerequisite and similarity relationships on different datasets. Learning Path Generation, in turn, improves the effectiveness of longer learning paths by dynamically incorporating similarity relationships between knowledge concepts.
4.1. Knowledge Structure Graph Generation
In this subsection, we present a concise overview of the KC Structure Graph Generation module. We begin by leveraging TextGrad (Yuksekgonul et al., 2025) to analyze the names of KCs, enabling more comprehensive retrieval of explanatory texts. Based on these texts, we design an EDU-GraphRAG framework that automatically constructs KC structure graphs, which serve as a crucial foundation for learning path recommendation.
Prompt for KC Explanation Generation LLM: (“You are an expert in the filed of education and are responsible for the detailed interpretation of knowledge concepts. The following are the knowledge concepts. Each knowledge concept is separated by ‘,’:knowledge concepts. Question: Use the knowledge in education to analyze the meaning of the serial number is begin to end. Knowledge concepts are: knowledge concepts[begin:end+1], as well as detailed analysis the relationship between each of these knowledge concepts and other knowledge concepts. Output Format: [knowledge concept]:[The analysis of knowledge concept.] Evaluation Instruction: Make the generated analysis smarter, more logical and accurate, more specific and discriminative rather than vague, and avoid ambiguity. Ensure that the analysis of knowledge concepts is correct.”)
4.1.1. Knowledge Concept Explanation Generation
To improve the reliability of KC explanations, we are inspired by the TextGrad framework (Yuksekgonul et al., 2025) to develop an iterative generation–evaluation– refinement procedure tailored to our task. We formulate the KC explanation refinement as a multi-turn optimization process. The initial explanation $\mathcal{T}^{(0)}$ is generated using a predefined prompt $P_{\mathrm{gen}}$. At each iteration $t$, the explanation $\mathcal{T}^{(t)}$ is evaluated by a secondary LLM with evaluation prompt $P_{\mathrm{eval}}$, producing feedback $\nabla_{\mathrm{LLM}}^{(t)}$ via an additional feedback prompt $P_{\mathrm{feedback}}$. The explanation is then refined with a rewriting prompt $P_{\mathrm{rewrite}}$ as:
(3) $\mathcal{T}^{(t+1)}=\mathrm{LLM}\left(P_{\mathrm{rewrite}},\mathcal{T}^{(t)},\nabla_{\mathrm{LLM}}^{(t)}\right),$
(4) $\nabla_{\mathrm{LLM}}^{(t)}=\mathrm{LLM}\left(P_{\mathrm{feedback}},\mathrm{LLM}\left(P_{\mathrm{eval}},c,\mathcal{T}^{(t)}\right)\right).$
This iterative process continues until convergence or a predefined maximum step $T$ is reached.
In this process, the knowledge concept name serves as the input. Through prompt engineering, the LLM is guided to act as a domain expert in education and generate an initial explanation. This explanation is then evaluated by another LLM to identify potential flaws, and iteratively refined based on the feedback until it reaches the desired level of accuracy and reliability. The finalized explanation is designated as the source document $D_{c}$ for the knowledge concept, to be consumed in the next stage of our framework. For instance, the objective function for the KC analysis is shown in the Prompt for KC Explanation Generation prompt box.
4.1.2. EDU-GraphRAG
Given a collection of $N$ concept-level explanation documents $\{D_{c_{1}},D_{c_{2}},\dots,D_{c_{N}}\}$, where each $D_{c_{i}}$ provides a contextualized explanation of a distinct knowledge concept $c_{i}$, we construct a unified document $D_{c}$ by concatenating all these explanations in sequence. Next, the document $D_{c}$ is segmented into a sequence of text chunks using a parameterized segmentation function $\sigma(D_{c};\theta)=\{\sigma_{1}(D_{c};\theta),\dots,\sigma_{n}(D_{c};\theta)\}$. For each chunk, an LLM abstractively extracts entities and relationships, capturing both explicit and implicit semantics in the text. These extracted components are passed to a graph constructor, which aggregates and deduplicates them into coherent local subgraphs. The global document-level graph index is obtained by unifying these subgraphs:
(5) $\phi(D_{c})=\bigcup_{i=1}^{n}\mathcal{G}\left(\mathcal{E}(\sigma_{i}(D_{c};\theta)),\mathcal{R}(\sigma_{i}(D_{c};\theta))\right).$
At query time, given a query $Q$, EDU-GraphRAG conditions an LLM on the structured graph index $\phi(D_{c})$ to generate the final output $A_{\mathcal{G}}$, which in our case corresponds to a constructed knowledge structure graph $\mathcal{G}$.
(6) $A_{\mathcal{G}}=\mathrm{LLM}(Q,\phi(D_{c})),$
where $A_{\mathcal{G}}=(\mathcal{C},\mathcal{S})$ denotes the induced knowledge structure graph comprising prerequisite and similarity relations among knowledge concepts, with $\mathcal{C}$ as the set of knowledge concepts and $\mathcal{S}$ as the set of semantic edges.
4.2. Learning Path Generation
In this subsection, we introduce the Discrimination Learning-driven Reinforcement Learning (DLRL) module, which dynamically generates personalized learning paths while addressing learning stagnation. To model the learner’s evolving knowledge state, we adopt the Difficulty Matching Knowledge Tracing (DIMKT) (Shen et al., 2022) as the environment. The decision-making process is guided by a multi-agent framework. Specifically, Prerequisite Agent (P-Agent) and Similarity Agent (S-Agent) jointly track learning progress and select the next KC based on either prerequisite or similarity relationships, and Difficulty Agent (D-Agent) selects exercises matched to the learner’s current ability level, ensuring a smooth progression in task difficulty.
4.2.1. Knowledge Tracing
The dynamic changes in a learner’s knowledge mastery level during the learning process serve as a reflection of their learning outcomes. By leveraging these insights to devise subsequent recommendation strategies, it is possible to provide a learning path that aligns with the learner’s abilities and progress. DIMKT is particularly advantageous as it can predict the impact of exercise difficulty on learners, thereby enhancing the accuracy of recommending exercises based on their difficulty level.
4.2.2. Prerequisite Agent
To ensure that the generated learning path follows a pedagogically appropriate order for the learner, we develop the P-agent to select knowledge concepts sequentially based on their prerequisite relationships.
a) State encoder: At each step, the predicted knowledge state is based on the learner’s learning objectives and historical learning records. Thus, at the end of the t1t-1 step, the state of the P-agent includes the learner’s knowledge state $\mathbf{h}_{t-1}$ and learning objectives $\mathbf{G}$. Specifically, we use a one-hot encoding scheme to represent the learning objective $\mathbf{G}=\{0,1\}^{M}$, where $M$ is the total number of knowledge concepts. The index corresponding to the learning objective is set to $1$, while all other indices are set to 0. At step $t$, the state of the P-agent $\mathbf{s}_{t}$ is represented as follows:
(7) $\mathbf{s}_{t}=\mathbf{h}_{t-1}\oplus\mathbf{G}.$